# Training Strategies for Isolated Sign Language Recognition

### v1

@@ Abstract
This paper introduces a comprehensive model training pipeline for Isolated Sign Language Recognition (ISLR) designed to accommodate the distinctive characteristics and constraints of the Sign Language (SL) domain. The constructed pipeline incorporates carefully selected image and video augmentations to tackle the challenges of low data quality and varying sign speeds. Including an additional regression head combined with IoU-balanced classification loss enhances the model’s awareness of the gesture and simplifies capturing temporal information. Extensive experiments demonstrate that the developed training pipeline easily adapts to different datasets and architectures. Additionally, the ablation study shows that each proposed component expands the potential to consider ISLR task specifics. The presented strategies improve recognition performance on a broad set of ISLR benchmarks. Moreover, we achieved a state-of-the-art result on the WLASL and Slovo benchmarks with 1.63% and 14.12% improvements compared to the previous best solution, respectively.
@@ 1 Introduction
Sign Languages (SL) are the primary means of communication for many deaf and hard of hearing individuals [^1] According to data from the All-Russian Society of the Deaf (VOG), there were more than 150 thousand native speakers of Russian Sign Language (RSL) in 2019.. Such languages do not replicate spoken speech but exhibit their own lexicon and unique grammatical rules. Due to the language gap, deaf and hard of hearing people may experience heightened prejudice in finding employment, pursuing academic education, or accessing medical services. Furthermore, learning sign language is challenging due to the limited number of teachers and native speakers who can help with communication practice.
Figure 1: The process of receiving IoU scores. (a) Consider the initial video, where the sign is between frames 2 and 5. (b) Random shift the sign’s boundaries by one frame to enhance model robustness (see Section 3.1 for details). (c) Collect frames by sliding a window of size 3 across the video. (d) Calculate IoU scores by dividing the number of sign frames in the window by the window size and adjust the classification scores (see Equation 2).
Sign language understanding is covered by three primary tasks: ISLR (Isolated Sign Language Recognition), CSLR (Continuous Sign Language Recognition), and SLT (Sign Language Translation). We begin the research of training strategies for SL problems with ISLR, which represents the video classification into individual signs. One of the beneficial applications of this task is to create an automatic sign language trainer, where users are shown a video example of a sign, and the recognition system assesses the quality of human gesturing. This advancement could make the learning of sign languages more accessible.
The development of automatic sign language recognition systems is becoming widespread [18, 10] to facilitate communication between deaf and hearing individuals. SLR is a complex challenge due to the impact of hand movement, location, orientation, shape, and facial expressions on sign meaning [34]. The SLR system should function in real-world settings (schools, hospitals, or train stations). This limitation introduces several hardships, including variations in sign display speeds, low video quality and resolution, diverse backgrounds, and varying lighting conditions. These factors adversely affect the system, which must operate with real-time response and achieve high quality since an error can have significant consequences. Besides real-world limitations, SLR is notably challenging due to the domain specifics, including the lack of available data and the direct temporal dependencies in videos.
@@
The basic approach to the ISLR task involves a classification neural network designed to process RGB video data. The model takes the sequence of frames sampled from the video as input and predicts a text label corresponding to the sign depicted in the video. The majority of current methodologies within this field provide either a solution at the model level [7, 45, 15], involving alterations to the model architecture, or at the data level [47], focusing on expanding the training data. In this work, we emphasize selecting an effective training strategy that considers the specificity of real-world SLR. Such a strategy holds the potential for substantial improvement without drastic alterations to the model or the dataset.
To address the domain-specific and real-world limitations, we propose a set of modifications for the basic approach, significantly enhancing the model quality across different datasets and architectures, including transformers [26, 38] and convolution neural networks (CNNs) [5]: (1) video-level data augmentations to simulate the different gesturing speeds (shown in Figure 3); (2) image-level data augmentations to reproduce low video quality; (3) an auxiliary sign boundary regression head to direct the network to focus more on the frames containing signs; and (4) 1D IoU-balanced CrossEntropy (IoU-balanced CE) loss to enhance the model’s capability to understand signs (see Figure 1). Considering the necessity for real-time response, we focus on modifying strategies solely based on RGB data without additional modalities like hand keypoints [30] and depth [21, 30] information.
@@
  • The versatile and scalable to various languages training pipeline for the ISLR models was created by considering the central of the task specifics.
  • The novel large-scale dataset, SlovoExt  which combines the Slovo [23] dataset with the newly assembled Russian Isolated Sign Language Dataset, detailed in Section 4. The code and pre-trained models are released to open-source [^2] https://github.com/ai-forever/TrainingStrategiesISLR to facilitate further research in the field. The SlovoExt dataset will be open-sourced as a part of an enormous RSL dataset in our future work.
  • Extensive experiments show that in comparison to the basic solution, MViTv2-S trained using the proposed pipeline is capable of achieving an increase of 6.54%, 3.93%, 10.12%, and 5.76% in top-1 accuracy on WLASL, AUTSL, Slovo and SlovoExt datasets, respectively. Table 3 shows the results of the ablation study of three proposed component blocks.
  • The state-of-the-art performance on two SLR datasets: (1) an 1.63% improvement relative to NLA-SLR [48] model on the widely-used American Sign Language benchmark WLASL [25] with 62.89 top-1 accuracy; (2) a 14.12% advancement relative to solution on the Russian Sign Language dataset Slovo with 78.21% mean accuracy. Such results are achieved by utilizing the presented pipeline with several additionals described in Section 5.4.
@@ 2 Related Work
Figure 3: Applying video augmentations to untrimmed videos. Identical frames are highlighted in the same color, and sign boundaries are outlined with a dashed line. Grey boundaries indicate duplicates of the last frame. a) speed up the video 2 times by removing every second frame; b) slow down the video 2 times by duplicating every second frame; c) random frames drop remains 80% of the total video length; d) random frames duplication increases the total video length by 20%; e) random boundary shift is applied, e.g., with shift of (1, -1), which means one frame is added on the left and one is removed on the right.
In this section, we present a summary of the SLR training strategies utilized within the proposed pipeline. We observe the widespread SLR studies and additionally consider action recognition and temporal action localization tasks.
@@
Sign Language Recognition Training Heuristics. Most works in the SLR field implemented model architecture-level alterations to improve performance. In the [14], self-mutual distillation was employed to enable the model to learn temporal and spatial features. In the[15], authors used correlation and identification modules to capture hands and face trajectories more effectively across consecutive frames. Other researchers employed data-based approaches to boost their model. The authors of [47] utilized datasets of various sign languages for training to address the issue of insufficient data. Some approaches incorporated not only RGB but also pose information [4], hand and body keypoints [22, 7, 45], and mouthing cues [33], combining data- and model-level modifications. This paper focuses on developing a training pipeline to be applied to any ISLR model that takes RGB videos as input and outputs classification scores.
@@
Image & Video Augmentations. Numerous studies focusing on SLR employed random crop [6, 7, 31, 1, 35] and horizontal flip [14, 31, 1, 35] as standard image augmentations. However, these studies should address handling non-mirrored signs, which should be kept the same since altering them could change their meaning. [44] utilized strong data augmentation techniques such as geometric and color space transformation to enhance the model’s robustness to data perturbations. We do not include such modifications due to inconsistency with real-life scenarios.
@@
Video augmentations are widely utilized in the action recognition task to enhance the model’s capacity to capture temporal dependencies. However, actions in this task (e.g., walking or running) are frequently repetitive and can be correctly classified at any time. On the contrary, the order of movements is essential in the SLR task. Such specific is the main reason for the impossibility of applying some of the proposed transformations in [12], where authors augmented data with video reversing, frame mixing, and temporal extension of CutMix [41]. In the [35], the authors proposed a video augmentation method where part of the video frames corresponding to one gesture are removed, added, or replaced with frames depicting a different gesture within the same video. This technique allowed authors to address the data deficiency and leverage the contextual information within the sign sequence. In this paper, we only remove and add frames (“speed up" and “slow down" in Section 3.1) to speed adjustments because replacing is unsuitable for the ISLR task, given only individual signs. [1] used different frame sampling rates to capture spatial and temporal information separately. We adopt a similar approach with randomness (“random add" and “random drop" in Section 3.1) to simulate real-life SLR cases. Such video augmentations, shown in Figure 3, enable sign recognition regardless of display speed and address data insufficiency in the SLR task.
@@
Auxiliary Regression Task. Localizing the action boundaries via regression loss can be beneficial when solving the action recognition task. It allowed the authors from [42] to achieve single-stage anchor-free temporal action localization. [46] also divided the task of temporal action localization into classification and action boundary regression and applied L1 loss for the second. As was stated in [27], supplementary tasks can boost the performance of the main task by pushing the backbone toward learning robust and generalized representations. Considering this concept, we incorporate an auxiliary sign boundary regression task by adding the regression head optimized by Huber loss [16] into the training pipeline. It assists the model in understanding the temporal position of the sign within the video, allowing it to focus more on the frames containing the sign.
@@
Intersection-over-Union Loss. Extracting information about an object’s spatial and temporal boundaries proves beneficial in both action recognition [40] and sign language recognition tasks [19, 33]. Techniques like IoU-loss [19, 29, 39] can aid this process. The authors [39] proposed employing the IoU-balanced CE loss to address the problem of the classification loss independence of localization predictions. [29] employed the 1D IoU loss in the action detection task to localize the video segment containing the action. [19] applied the Generalized IoU loss for sign language recognition: the model processed images depicting signs and calculated IoU between bounding boxes of hands within the frame. In the proposed pipeline, we combine the concepts of 1D IoU and the IoU-balanced classification losses to devise the classification IoU loss tailored for the video domain and the particularities of the ISLR task. When computing the classification score, this approach allows us to consider the relative localization of the sign (see Figure 1 for details).
@@ 3 Training Strategies
Due to the specificity of the sign language domain, the design of a training pipeline requires particular approaches that consider both the data and the model requirements. Figure 2 shows the proposed strategies divided into three components: (1) video-level augmentations, (2) image-level augmentations, and (3) additional losses. Note that these techniques can be utilized with any SL dataset and architecture except for several subcomponents mentioned in Section 3.2.
3.1 Video Augmentations
@@
Speed Up & Slow Down. In real life, signs are shown at various speeds, which poses challenges in crowding a dataset with a real-like distribution due to a limited number of sign language speakers. Therefore, during model training, we artificially speed up or slow down the video recordings without overstepping the limits of gesturing speed encountered in the real world. We sample every N-th frame from the video sequence to speed up the video by a factor of N, where N is an integer (see Figure 3a). The additional frames from the original video are pre-sampled to the right to maintain the clip length after acceleration. We repeat each frame N times to slow down the video (see Figure 3b), pre-dropping the frames on the right so that after the subsequent addition, the clip length remains the initial.
Random Add & Random Drop. In some real-life scenarios, gesturing speed is non-uniform because of the difficulty or simplicity of some gesture parts or due to external distractions during the showing. We partially change the speed of the videos to make the model more resistant to such uncommon cases. Figure 3c illustrates the involvement of a random drop of frames from the video sequence to speed up the video partially (see Algorithm 1). As in the case of uniform video speed up, we retain the same clip length using the identical process. Similarly, a random add of frames is applied to the video for a partial slow down (see Figure 3d). We are maintaining the video length in the same way as in a uniform slow down.
Random Boundary Shift. During inference, an SLR system generally captures frames sequentially with a window of a certain length. So, too short or too long signs may not fit into the window properly. Therefore, to ensure valid recognition of such signs, we augment the data by shifting the sign boundaries by a random value, removing part of the signs, or adding frames without the signs (see Figure 1 and Figure 3e for illustrations).
@@ 3.2 Image Augmentations
Image Quality. We imitate artifacts in video recordings caused by low-resolution capturing, a defect in the video camera, or problems with video transmission over the network by compressing and downscaling images and adding random noise and sharpness.
Horizontal Flip. Most signs’ meanings remain unchanged after horizontal mirroring. Hence, horizontal flip transformation facilitates the equal processing of signs shown by different hands. However, some signs are not invariant to the horizontal flip (e.g., RSL gestures such as “left”, “right”, “heart”, and “liver” stop transferring their meanings if shown with the inappropriate hand). In the experiments, this augmentation only affects mirrored signs (more details in Section 5.1).
General. In addition to domain-specific augmentations, we utilize general ones to diversify the dataset: color jittering to provide heterogeneity over the color context of the frame, and CutMix [41] and MixUp [43] to induce the model to learn more generalizable features.
@@ 3.3 Additional Losses
The following domain-specific training techniques incorporate domain knowledge of the sign boundaries into the model by scaling the classification loss and solving an additional task to enhance the main task’s solution accuracy. Such modifications are not model-level, as they can be integrated into either architecture.
@@
Classification Loss Scaling. Conventional classification losses like cross-entropy loss have the following limitation: they do not consider information about the localization of the sign in the video, which can cause irrelevant gradients to optimize. We use a IoU-balanced CE loss to incorporate information about the location of the sign relative to the window. To calculate IoU, the length of the intersection of the sign with the sampled window is divided by the size of the window:
$IoU=\frac{\min(w_{end},s_{end})-\max(w_{start},s_{start})}{w_{end}-w_{start}+1},$ (1)
where $s_{start}$, $w_{start}$ are the first frames of the sign and the sampled window, and $s_{end}$, $w_{end}$ are the last frames, respectively (for illustration, see Figure 1 and Figure 2). The overall classification loss is the weighted combination of classification scores and IoU coefficients, denoted as:
$\mathcal{L}_{cls}=CrossEntropy(iou_{i}\ast p_{i},\hat{p}_{i}),$ (2)
where $p_{i}$ and $\hat{p}_{i}$ are the predicted classification score and its corresponding ground-truth label, and $iou_{i}$ is the IoU score of the video.

### v2

@@ Abstract
Accurate recognition and interpretation of sign language are crucial for enhancing communication accessibility for deaf and hard of hearing individuals. However, current approaches of Isolated Sign Language Recognition (ISLR) often face challenges such as low data quality and variability in gesturing speed. This paper introduces a comprehensive model training pipeline for ISLR designed to accommodate the distinctive characteristics and constraints of the Sign Language (SL) domain. The constructed pipeline incorporates carefully selected image and video augmentations to tackle the challenges of low data quality and varying sign speeds. Including an additional regression head combined with IoU-balanced classification loss enhances the model’s awareness of the gesture and simplifies capturing temporal information. Extensive experiments demonstrate that the developed training pipeline easily adapts to different datasets and architectures. Additionally, the ablation study shows that each proposed component expands the potential to consider ISLR task specifics. The presented strategies enhance recognition performance across various ISLR benchmarks and achieve state-of-the-art results on the WLASL and Slovo datasets.
@@ 1 Introduction
Sign Languages are the primary means of communication for many deaf and hard of hearing individuals. According to data from the All-Russian Society of the Deaf (VOG [^1] https://voginfo.ru/all-russian-society-of-the-deaf/), there were more than 150,000 native speakers of Russian Sign Language (RSL) in 2019. Such languages do not replicate spoken language but possess their own lexicon and unique grammatical rules. Due to the language gap, deaf and hard of hearing people may experience prejudice in finding employment, pursuing academic education, or accessing medical services. Furthermore, learning Sign Language (SL) is challenging due to the limited number of teachers and native speakers.
However, developing robust Sign Language Recognition (SLR) systems remains challenging. Variations in gesturing speed, complex spatial-temporal features, and real-world capturing conditions (e.g., diverse backgrounds or lighting) can notably affect recognition accuracy. Furthermore, large-scale, high-quality datasets are still limited.
Figure 1: The process of receiving IoU scores. (a) Consider the initial video, where the sign is between frames 2 and 5. (b) Randomly shift the sign‘s boundaries by one frame to enhance model robustness (see Section 3.1 for details). (c) Collect frames by sliding a window of size 3 across the video (window size of 3 chosen only for illustration). (d) Calculate IoU scores by dividing the number of sign frames in the window by the window size and adjust the classification scores. Note that the window size of 3 in this figure was selected purely as an example and was not used in our experiments.
Sign language understanding is covered by three primary tasks: ISLR (Isolated Sign Language Recognition), CSLR (Continuous Sign Language Recognition), and SLT (Sign Language Translation). This research focuses on ISLR, where videos are classified into individual sign categories. One of the beneficial applications of this task is to create an automatic SL trainer, where users are shown a video example of a sign, and the recognition system assesses the quality of human gesturing. This advancement could make the learning of SL more accessible.
The development of automatic SLR systems is becoming widespread [16, 8] to facilitate communication between deaf and hearing individuals. SLR is a complex challenge due to the impact of hand movement, location, orientation, shape, and facial expressions on sign meaning [33]. The SLR system should function in real-world settings (schools, hospitals, or train stations). Its complexity is due to variations in sign display speeds, low video quality and resolution, diverse backgrounds, and varying lighting conditions. These factors adversely affect the system, which needs to operate with real-time response and maintain high quality to minimize errors. Besides real-world limitations, SLR faces domain-specific challenges such as limited data availability and complex temporal dependencies. Current methodologies often do not fully address these issues, limiting the development of models that generalize across diverse signing styles, varying execution speeds, and other conditions. To address these problems, we explore novel training strategies and propose a robust pipeline designed to handle the aforementioned challenges in the ISLR domain.
@@
The basic approach to the ISLR task involves a classification neural network, designed to process RGB video data. The model takes the sequence of frames sampled from the video as input and predicts a text label corresponding to the sign depicted in the video. Many existing methodologies within this field provide either a solution at the model level [6, 44, 13], involving alterations to the model architecture, or at the data level [46], focusing on expanding the training data. Due to unaddressed domain-specific challenges, these approaches may not fully leverage the models’ potential for ISLR. Our work introduces a training pipeline designed for real-world SLR. It boosts performance without altering the underlying model architecture or dataset. To address both domain-specific challenges and practical deployment constraints, we apply a series of modifications to the basic approach. These enhancements improve accuracy across multiple datasets and model types, including transformers [24, 37] and convolutional neural networks (CNNs) [4]: (1) video-level data augmentations to simulate the different gesturing speeds (shown in Figure 3); (2) image-level data augmentations to reproduce low video quality; (3) an auxiliary sign boundary regression head to direct the network to focus more on the frames containing signs; and (4) 1D Intersection-over-Union-balanced CrossEntropy (IoU-balanced CE) loss to enhance the model’s capability to understand signs (see Figure 1). Considering the necessity for real-time response, we focus on modifying strategies solely based on RGB data without additional modalities like hand keypoints [29] and depth [19, 29] information.
@@
  • We propose a versatile and scalable training pipeline for ISLR models that considers domain-specific constraints, including data quality, varying sign execution speeds, and sign boundary awareness.
  • We present a novel large-scale dataset, SlovoExt, which combines the Slovo [21] dataset with our newly assembled Russian Isolated Sign Language Dataset; we also release [^2] https://github.com/ai-forever/TrainingStrategiesISLR the code and pre-trained models to facilitate further research. The SlovoExt dataset will be available as part of a larger RSL dataset in future work.
  • We illustrate how the proposed training strategies enable superior performance compared to existing solutions, bridging the gap between data-level and model-level improvements in SLR tasks.
@@ 2 Related Work
Sign Language Recognition Training Heuristics. Many works in the SLR field have implemented model architecture-level alterations to improve performance. In [12], self-mutual distillation was employed to enable the model to learn temporal and spatial features. In [13], authors used correlation and identification modules to capture hands and face trajectories more effectively across consecutive frames. Other researchers have employed data-based approaches to boost their model. The authors of [46] utilized datasets of various sign languages for training to address the issue of insufficient data. Some approaches have incorporated not only RGB but also pose information [3], hand and body keypoints [20, 6, 44], and mouthing cues [32], combining data- and model-level modifications. This paper focuses on developing a training pipeline to be applied to any ISLR model that takes RGB videos as input and outputs classification scores.
@@
Image & Video Augmentations. Many studies focusing on SLR have employed random crop [5, 6, 30, 1, 34] and horizontal flip [12, 30, 1, 34] as standard image augmentations. However, these studies often do not address how to handle non-mirrored signs, which should be kept the same since altering them could change their meaning. Zhou et al. [43] utilized strong data augmentation techniques, such as geometric and color space transformations, to enhance the model’s robustness to data perturbations.
@@
Video augmentations have proven widely effective in the action recognition task to enhance the model’s capacity to capture temporal dependencies. However, actions in this task (e.g., walking or running) are frequently repetitive and can be correctly classified at any time. On the contrary, the order of movements is essential in the SLR task. The specificity is the main reason for the inability to apply some of the transformation proposed in [10], where authors augmented data with video reversing, frame mixing, and temporal extension of CutMix [40]. In [34], the authors proposed a video augmentation method where part of the video frames corresponding to one gesture is removed, added, or replaced with frames depicting a different gesture within the same video. This technique allowed the authors to address the data deficiency and leverage the contextual information within the sign sequence. In this paper, we only remove and add frames (“speed up" and “slow down" in Section 3.1) to speed adjustments because replacing is unsuitable for the ISLR task, given that only isolated signs are present. Ahn et al. [1] used different frame sampling rates to capture spatial and temporal information separately. We adopt a similar approach with randomness (“random add" and “random drop" in Section 3.1) to simulate real-life SLR cases. These video augmentations, shown in Figure 3, enable sign recognition regardless of display speed and address data insufficiency in the SLR task.
@@
Auxiliary Regression Task. Localizing the action boundaries via regression loss has demonstrated benefits in solving action recognition tasks. This approach enabled Zhang et al. [41] to achieve single-stage anchor-free temporal action localization. Similarly, the authors of [45] divided the temporal action localization task into classification and action boundary regression, applying L1 loss for the latter. As was stated in [25], supplementary tasks can boost the performance of the main task by pushing the backbone toward learning robust and generalized representations. Considering this concept, we incorporate an auxiliary sign boundary regression task by adding a regression head optimized by Huber loss [14] to the training pipeline, as it has demonstrated better model convergence in the experiments provided below (see Table 5). It assists the model in understanding the temporal position of the sign within the video, allowing it to focus more on the frames containing the sign.
@@
IoU-balanced CE Loss. Extracting information about an object’s spatial and temporal boundaries proves beneficial in both action recognition [39] and SLR tasks [17, 32]. Techniques like IoU-loss [17, 27, 38] can aid this process. Wu et al. [38] proposed employing the IoU-balanced CE loss to address the independence between classification and localization predictions. Liu et al. [27] employed a 1D IoU loss in the action detection task to localize relevant segments within the video. In [17], the Generalized IoU loss was employed for SLR by calculating IoU between the bounding boxes of hands in each frame. In the proposed pipeline, we combine the concepts of 1D IoU and the IoU-balanced classification losses to devise the classification IoU loss tailored for the video domain and the particularities of the ISLR task. When computing the classification score, this approach allows us to consider the relative localization of the sign (see Figure 1 for details).
@@ 3 Training Strategies
To address the existing gap in SLR methodologies, designing an effective training pipeline requires tailored approaches that account for both data characteristics and model requirements. Figure 2 shows the proposed strategies divided into three components: (1) video-level augmentations, (2) image-level augmentations, and (3) additional losses. Note that these techniques do not depend on specific SL datasets or model architectures, except for certain subcomponents detailed in Section 3.2.
3.1 Video Augmentation
@@
Speed Up & Slow Down. In real life, signs can appear at various speeds, so we artificially speed up or slow down videos to match the real-world gesturing speed distribution. We sample every N-th frame to speed up by a factor of N (see Figure 3a), pre-sampling additional frames on the right to maintain the same clip length. Conversely, we repeat each frame N times to slow down the video (see Figure 3b), pre-dropping frames so that the final length remains unchanged.
Figure 3: Applying video augmentations to untrimmed videos. Identical frames are highlighted in the same color, and sign boundaries are outlined with a dashed line. Grey boundaries indicate duplicates of the last frame. a) speed up the video 2 times by removing every second frame; b) slow down the video 2 times by duplicating every second frame; c) random frames drop remains 80% of the total video length; d) random frames duplication increases the total video length by 20%; e) random boundary shift is applied, e.g., with shift of (1, -1), which means one frame is added on the left and one is removed on the right. These values were chosen for ease of demonstration.
Random Add & Random Drop. In some real-life scenarios, gesturing speed is non-uniform because of the difficulty or simplicity of some gesture parts or due to external distractions during the hand movement. We partially change the speed of the videos to make the model more resistant to such uncommon cases. Figure 3c illustrates the process of randomly dropping frames to speed up the video partially (see Algorithm 1). As in the case of uniform video speed up, we retain the same clip length using the identical process. Similarly, a random add of frames is applied to the video for a partial slow down (see Figure 3d). We are maintaining the video length in the same way as in a uniform slow down.
Random Boundary Shift. During inference, an SLR system generally captures frames sequentially using a window of a specific length. Therefore, signs that are too short or too long may not fit into the window properly. To ensure accurate recognition, we augment the data by randomly shifting sign boundaries, removing part of the videos, or adding frames without signs (see Figure 1 and Figure 3e for illustrations).
@@ 3.2 Image Augmentation
Image Quality. We imitate artifacts in video recordings caused by low-resolution capturing, due to defects in the video camera, or problems with video transmission over the network by compressing and downscaling images and adding random noise and sharpness.
Horizontal Flip. Most meanings of the signs remain unchanged after horizontal mirroring. Hence, horizontal flip transformation facilitates the equal processing of signs shown by different hands. However, some signs are not invariant to the horizontal flip (e.g., RSL gestures such as “left”, “right”, “heart”, and “liver” stop transferring their meanings if shown with the inappropriate hand). In the experiments, this augmentation only affects mirrored signs (more details in Section 5.1).
General. In addition to domain-specific augmentations, we utilize general ones to diversify the dataset: color jittering to provide heterogeneity over the color context, and CutMix [40] and MixUp [42] to induce the model to learn more generalizable features.
@@ 3.3 Additional Losses
The following domain-specific training techniques incorporate domain knowledge of the sign boundaries into the model by scaling the classification loss and solving an additional task to enhance the main task’s solution accuracy. These modifications are not model-level, as they can be integrated into either architecture.
@@
Classification Loss Scaling. Conventional classification losses, like cross-entropy, have the limitation to not consider information about the localization of the sign in the video, which can cause irrelevant gradients to optimize. We use a IoU-balanced CE loss to incorporate information about the location of the sign relative to the window. To calculate IoU scores, the length of the intersection of the sign with the sampled window is divided by the size of the window:
$IoUscore=\frac{\min(w_{end},s_{end})-\max(w_{start},s_{start})}{w_{end}-w_{start}+1},$ (1)
where $s_{start}$, $w_{start}$ are the first frames of the sign and the sampled window, and $s_{end}$, $w_{end}$ are the last ones, respectively; $min(a,b)$ and $max(a,b)$ functions denote the minimum and maximum of two values, respectively (for illustration, see Figure 1 and Figure 2). The overall classification loss is the combination of classification and IoU scores.